OpenStudy (anonymous):

Convert the polar equation r^2(1+2sin^2theta)=4 to an equation with Cartesian (rectangular) coordinates

OpenStudy (dumbcow):

Note: x = r*cos(theta) y = r*sin(theta) x^2 +y^2 = r^2 --> r^2 + 2r^2 sin^2 = 4 --> (x^2 +y^2) +2y^2 = 4 -->x^2 +3y^2 = 4 looks like an ellipse

OpenStudy (anonymous):

solution from book said x^2 + (y+1)^2=5

OpenStudy (anonymous):

I think dumbcow is right...

OpenStudy (anonymous):

according to your solution your center is at (0,-1) but mathematica has it at (0, 0)

OpenStudy (anonymous):

yea i trust dumbcow

OpenStudy (anonymous):

thanks guys

OpenStudy (dumbcow):

welcome, the solution doesn't match are you sure the question is typed correctly

OpenStudy (anonymous):

the solution is wack lots of mistakes in it lol

OpenStudy (dumbcow):

haha ok well i double checked on wolfram and i am correct

OpenStudy (anonymous):

one more thing... x^2 + (y+1)^2=5 is actually a cirlce with radius sqrt5

OpenStudy (anonymous):

does wolfram teaches how to do it? or you just look at the graph and can tell right away

OpenStudy (dumbcow):

no not that i know of i just compared the graphs

OpenStudy (anonymous):

thanks have a good night!